How many liters of 20% alcohol solution must be mixed with 60 liters of 70% solution to get a 60% solution?
Answer by mananth(16946)

--------percent ---------------- quantity
Alcohol I 70 ---------------- 60 L
Alcohol II 20 ---------------- x L
Mixture 60 ---------------- 60 + x L

70*60+20x=60(60+x)
4200 + 20 x =3600+60x
20 x + -60x=3600-4200
-40 x = -600
/ -40
x = 15
15 L of Alcohol II